# Energy-Efficient Balanced Flow Control Achieved through Optimization of Synthetic Jet Placement Based on Deep Reinforcement Learning

### v1

@@ 2.5 DRL-Enhanced Active Flow Control
  • Agent : The neural network of the agent is parameterized by $\theta$ (including weights and biases), and through training, this network learns the optimal policy $\pi_{\theta}(a_{t}\mid s_{t})$, which represents the probability distribution of actions $a_{t}$ given a state $s_{t}$. The objective is to maximize the expected cumulative reward $R_{t}=\sum_{k>t}\gamma^{(k-t)}r_{k}$, where $\gamma\in(0,1]$ is the discount factor. PPO is an episode-based actor-critic algorithm that includes both an actor network $\pi_{\theta}$ and a critic network $V(s_{t})$. In the PPO algorithm, the actor network is responsible for learning and outputting the policy $\pi_{\theta}(a_{t}\mid s_{t})$ for selecting actions in different states to maximize the cumulative reward $R_{t}$. The critic network $V(s_{t})$ evaluates the expected cumulative reward for each state $s_{t}$ and provides feedback and guidance for updating the policy $\pi_{\theta}$ of the actor network. Both the actor and critic networks are comprised of two fully connected layers, each with 512 units. The Adam optimizer is used for updating the networks, with a fixed learning rate of 0.001. In reinforcement learning, agent selection depends on the state space, action space, exploration-exploitation trade-off, computational resources, and objectives. For high-dimensional continuous spaces like fluid mechanics, PPO (Rabault et al. (2019, 2020); Tang et al. (2020a)), SAC (Wang et al. (2022a); Wang & Xu (2024d); Xia et al. (2024)), TD3 (Fan et al. (2020)) and TQC (Xia et al. (2024)) are ideal, with PPO excelling in real-time flow control due to its exploration balance and efficiency.
@@
  • Environment : the environment interacting with the agent is instantiated using CFD. We utilized the gym library interface to integrate OpenFOAM into the DRL framework. The CFD simulations are performed using the open-source software platform OpenFOAM® version 8. Figure 7 illustrates the integration of CFD computations as a DRL training environment and demonstrates the benefits of accelerating the training process by using multiple parallel CFD environments. This parallelization strategy is inspired by the work of Rabault & Kuhnle, and also incorporates insights from Wang & Xu. Research by Rabault & Kuhnle indicates that acceleration can be achieved by parallelizing both the numerical simulations and the DRL algorithm. In their study, CFD computations utilized the finite element method implemented within the FEniCS framework. Wang & Xu demonstrated that different CFD computation platforms can significantly affect parallel acceleration. They investigated various hybrid parallelization configurations based on the OpenFOAM environment and proposed effective parallelization strategies. This study involves two flow scenarios: flow around a square cylinder and flow around a circular cylinder, which were described in detail in the previous sections. The environment depends on the type of CFD solver, and there are various codes available for coupling numerical simulation environments with RL frameworks. For example, RL has been integrated with the FEniCS environment (Rabault et al. (2019)), the OpenFOAM environment (Wang et al. (2022a)), and the Nek5000 environment (Li & Zhang (2022)).
[…]
    $V_{\Gamma_{1},T_{i}}=V_{\Gamma_{1},T_{i}}+\beta\left(a-V_{\Gamma_{1},T_{i-1}}\right).$ (14)
  • State $s_{t}$ : The agent observes the state of the CFD environment by collecting physical field data from specific locations within the computational domain. Building on previous studies, such as those by Wang et al., who utilize velocity probes, and Rabault et al., who employ pressure probes to capture instantaneous pressure at key points, we draw inspiration from Li & Zhang to strategically place observation probes at locations within the flow field where fluctuations are most pronounced. These probes are situated in regions of the wake characterized by the weakest stability and highest sensitivity, allowing for more effective flow control through the monitoring of time-averaged fluctuating pressure. This targeted placement optimizes the control strategy by focusing on the most critical areas of the flow, thereby enhancing both the efficiency and effectiveness of the flow control efforts. In addition to positioning probes within the wake region, we also place probes around the circumference of the cylinder to extract physical information from its surface, which is closely related to the lift and drag on the cylinder. For more details on the number and distribution of these probes, readers can refer to the studies by Paris et al.; Li & Zhang; Wang et al., who have conducted in-depth discussions on this topic.
  • Reward $r_{t}$ : The design of the reward function is based on the overall objective of flow control, making the selection of appropriate performance metrics crucial for the convergence and stability of reinforcement learning. In this study, we aim for flow control to reduce both drag and lift around the cylinder. Therefore, we construct the reward function based on the drag coefficient ($C_{D}$) and the lift coefficient ($C_{L}$). The specific reward function is defined as follows:
@@
    where, $C_{D,0}$ represents the baseline drag coefficient, serving as a reference to encourage positive rewards and facilitate convergence during training. The term $(C_{D})_{T_{i}}$ denotes the drag coefficient at time step $T_{i}$, with the objective of minimizing it relative to the baseline. The lift coefficient $(C_{L})_{T_{i}}$ is also included, with its absolute value penalized to mitigate instabilities caused by lift around the bluff body. The parameter $\omega$ is a weighting factor used to balance the trade-off between reducing drag and controlling lift. Given the directional nature of the flow under consideration, $\omega$ is typically set between 0.1, 0.2 and 1, reflecting the relative importance of minimizing drag over controlling lift fluctuations. In previous work, the lift coefficient penalty factor is often set to $0.1$ (Wang & Xu, 2024c), $0.2$ (Rabault et al., 2019; Tang et al., 2020b), or $1$ (Ren et al., 2021). In this study, we use a penalty factor of 0.2, consistent with Rabault et al..
  • Interaction: In DRL, the interaction between the agent and the environment forms the foundation of the learning process. At each time step $t$, the agent first gathers physical data regarding the current flow state $s_{t}$ from the CFD environment. Based on this information, the agent selects an action $a_{t}$ according to its control policy $\pi_{\theta}(a_{t}\mid s_{t})$. This selected action $a_{t}$ is then communicated to the CFD environment, which executes the action, transitions to the next state $s_{t+1}$, and provides an immediate reward $r_{t}$. The agent, in turn, receives both the reward $r_{t}$ and the updated physical state information $s_{t+1}$, using this feedback to refine its control policy $\pi_{\theta}$ in order to maximize cumulative future rewards $\sum_{t}\gamma^{t}r_{t}$, where $\gamma$ is the discount factor. This iterative cycle continues until the task is accomplished or a predefined termination condition is met.
    Through successive interactions, the agent incrementally optimizes its control policy $\pi_{\theta}$ via trial and error, ultimately mastering the dynamic adjustment of jet velocity $V_{\Gamma}$ in real-time to achieve optimal flow control performance. In this study, the total number of DRL training episodes is set to 3,000, with each episode consisting of 100 training timesteps. Each timestep corresponds to a non-dimensional time of 0.025, resulting in a maximum duration per episode of $T_{\text{max}}=2.5$. This duration spans approximately 8 to 10 vortex shedding cycles, during which 100 control actions are executed. Training the DRL agent over multiple vortex shedding cycles ensures that the agent fully captures the complex flow dynamics, stabilizes the learning process, and develops a robust and generalizable control strategy for effective flow management.
@@ 3 Results: flow control performance of a circular cylinder
In this section, we employ the flow around a cylinder as the DRL training environment, utilizing the control capabilities of a PPO agent to achieve active flow control through synthetic jets, with the primary objectives of drag reduction and vortex suppression. The focus of this study is to investigate how the positioning of synthetic jets affects flow control performance and the associated costs, with the aim of deriving qualitative insights through rigorous physical analysis. The cost of implementing flow control is quantified by the dimensionless mass flow rates of the synthetic jets. To explore the effects of jet positioning, we design five configurations involving a single pair of synthetic jets, two configurations involving two pairs of synthetic jets, and one additional configuration with a single pair of synthetic jets. Each scheme is carefully constructed to assess the influence of jet distribution on the overall flow control effectiveness and efficiency.
@@ 3.1.1 Scenario A: Control Performance of Jet Pair
Figure 8: Performance evaluation of different jets configurations for flow control around a cylinder. (a) Reward curves over training episodes for various jets positions. (b) Drag coefficient $C_{D}$ over non-dimensional time, comparing the baseline scenario with different jets configurations. (c) Lift coefficient $C_{L}$ over non-dimensional time, comparing the baseline scenario with different jets configurations.
@@
We compare the performance and effectiveness of flow control strategies obtained through DRL training with synthetic jets placed at different positions on the cylinder surface. These five synthetic jets configurations are designed based on the angular positions on the cylinder, ranging from 60° to 120°. In these configurations, the $J_{1}$ setup includes two jets, $J_{1}^{1}$ and $J_{1}^{2}$, with $J_{1}^{1}$ located at an angular position of 120°. The $J_{3}$ setup includes two jets, $J_{3}^{1}$ and $J_{3}^{2}$, with $J_{3}^{1}$ located at 90°. The $J_{5}$ setup also includes two jets, $J_{5}^{1}$ and $J_{5}^{2}$, with $J_{5}^{1}$ located at 60°.
As shown in figure 8, the reward function curves reflect how the performance of the DRL agent evolves as training progresses. In the early stages of training, the reward function curves often exhibit significant fluctuations, indicating the exploration and experimentation phase of the agent within the strategy space. As training continues, the agent gradually optimizes its strategy, leading to an upward trend in the reward function values, which suggests that the agent is effectively learning the flow control strategy. By comparing the learning curves for different jet positions, we observe that when the jets are positioned at $J_{1}$, the reward function grows most slowly and reaches the lowest final value, indicating that the flow control strategy at this position is more challenging to converge. In contrast, when the jets are positioned at $J_{5}$, the learning curve shows a slightly faster growth but still lags behind others, indicating a moderate difficulty in strategy convergence. Conversely, when the jets are positioned at $J_{2}$, $J_{3}$, and $J_{4}$, the DRL agent achieves rapid convergence within 1,000 episodes, demonstrating that the flow control strategies at these positions are easier to learn and optimize. Between 1,000 and 3,000 episodes, the reward function curves plateau, indicating that the DRL agent has successfully mastered a stable and effective flow control strategy. At this stage, the reward values remain stable, further confirming the effectiveness and reliability of the learned strategy in reducing drag and controlling lift.
The baseline $C_{D}$ represents the scenario without any active flow control and serves as a reference point. As shown in figure 8, all jet configurations achieve a reduction in $C_{D}$, with drag reduction of approximately 8% across the five jet positions. This indicates that the DRL-optimized jet actuation effectively reduces the drag experienced by the cylinder, with the curves stabilizing after the initial transient phase, suggesting that the control actions lead to a steady reduction in drag. For the same jet configurations, the baseline $C_{L}$ exhibits significant fluctuations, which are common in flows past bluff bodies such as cylinders. The DRL-controlled jets achieve a substantial reduction in these fluctuations, with the lift oscillations being nearly eliminated in all five jet positions, resulting in up to a 99% reduction in $C_{L}$ oscillations. The dramatic reduction in lift fluctuations indicates that the jets not only reduce drag but also stabilize the flow, preventing lift-induced instabilities that could otherwise affect the effectiveness of flow control. The stabilization of the drag and lift coefficients over time demonstrates that the learned strategies are both effective and stable. The DRL training curves illustrate the learning process of the agent, resulting in an optimized control strategy that effectively reduces both drag and lift coefficients. The significant drag reduction and near-complete suppression of lift fluctuations highlight the potential of DRL in achieving robust and efficient flow control in real-time applications.
The activation states of synthetic jets at various azimuthal angles, along with the corresponding controlled flow fields, are illustrated in figure 9. When the activation states of synthetic jets at different azimuthal angles are shown in the left column of figure 9, the pair of synthetic jets symmetric about the x-axis alternately engage in blowing or suction phases. With the jets positioned at $J_{1}$ at 120°, the velocity contours show the development of a large recirculation region on the cylinder’s rear side, despite minor oscillations in the wake. When the jets are positioned at $J_{2}$ at 105° and $J_{3}$ at 90°, vortex shedding in the wake is fully suppressed, indicating that jets at these locations can effectively delay separation, reduce the separation zone, and even eliminate it, thereby reducing drag. For $J_{4}$ at 75°, the wake is largely stabilized, though weak oscillations persist downstream. Glezer & Amitay (2002) used smoke flow visualization to show that a synthetic jet at an injection angle of $\gamma=60^{\circ}$ caused local deformation of the streaklines on the upper surface of the cylinder, shifting the separation point downstream and the front stagnation point below the x-axis. When the jet is set at $\gamma=60^{\circ}$, our reinforcement learning algorithm achieves similar control performance over the cylinder flow field as observed by Glezer & Amitay (2002), with periodic vortex shedding still evident in the controlled flow. The effectiveness of flow control varies significantly with jet positioning, with the least effective control observed at $J_{1}$ (120°) and $J_{5}$ (60°), where the jets can still delay separation.
@@ 3.1.2 Scenario A: Energy Consumption of Jet Pair
Figure 10: Energy consumption analysis of jet pairs positioned at different angular locations around a cylinder. The left panel illustrates the angular placement of jet pairs $J_{1}$ through $J_{5}$ around the cylinder, while the right panel shows the corresponding action values over non-dimensional time for each jet configuration.
A detailed analysis of the energy consumption associated with various jet configurations around a cylinder, along with their effectiveness in flow control, is presented in figure 10. The figure depicts the angular placement of five jet pairs ($J_{1}$ through $J_{5}$) along the cylinder surface, with $J_{1}$ located nearest to the forward stagnation point (0°) and $J_{5}$ positioned further downstream. These jet pairs are symmetrically distributed on both sides of the cylinder’s centerline, covering angular positions from 120° to 60°. The right side of the figure shows the action values, representing the intensity of jet actuation, plotted over non-dimensional time for each jet configuration, thereby providing insight into the energy consumption required to achieve the desired flow control.
The curves for each configuration initially exhibit a peak, representing the maximum energy required at the onset of jet actuation. Notably, the curves for $J_{5}$ and $J_{1}$ display the highest initial action values, indicating that these positions demand more energy to initiate flow control compared to other configurations. Following the initial transient phase, the action values stabilize, signifying that the control effort, and thus energy consumption, becomes more consistent over time. The action values for configurations $J_{2}$, $J_{3}$, and $J_{4}$ stabilize at lower levels than those for $J_{5}$ and $J_{1}$, suggesting that these configurations are more energy-efficient in maintaining flow control. The annotations on the plot highlight that energy consumption peaks around the 90° position, while it remains comparable at 75° and 105°. This observation implies that jets positioned at 60° and 120° (such as $J_{5}$ and $J_{1}$) are less energy-efficient, requiring more energy to achieve the same flow control effect as the more efficiently positioned jets at 90° and 105°. This insight highlights the critical importance of optimizing jet positioning to minimize energy consumption, a key consideration in designing efficient flow control systems that maximize aerodynamic performance while reducing power requirements around the cylinder.
@@ 3.2.1 Scenario B: Control Performance of Multiple Jets Control
Figure 11: Control performance evaluation of multiple jet configurations around a cylinder. (a) Reward curves over training episodes for various multi-jet configurations. (b) Drag coefficient $C_{D}$ over non-dimensional time. (c) Lift coefficient $C_{L}$ over non-dimensional time.
Based on the results of flow control using a single pair of synthetic jets, it was observed that the configurations with jets positioned at 60° and 120° ($J_{5}$ and $J_{1}$) consumed the most energy. Consequently, when employing multiple pairs of synthetic jets for flow control, we excluded the jets positioned at 60° and 120°, focusing instead on combinations involving $J_{2}$, $J_{3}$, and $J_{4}$. The configurations tested include two pairs of synthetic jets ($J_{2}$ and $J_{3}$; $J_{3}$ and $J_{4}$) and a configuration with three pairs ($J_{2}$, $J_{3}$, and $J_{4}$). To investigate the control performance and convergence speed of multiple synthetic jets combinations compared to single-pair synthetic jets, we analyze the control performance of single-pair synthetic jets ($J_{2}$, $J_{3}$, and $J_{4}$) as the control group.
The performance of multiple jet configurations used for flow control around a cylinder, evaluated through deep Reinforcement Learning training, is comprehensively analyzed in figure 11(a). The reward curves reflect the effectiveness of the DRL agent in learning flow control strategies using different combinations of synthetic jets configurations. Across various configurations, the reward function exhibits a stepwise increase during the initial training phase. As training progresses, the reward values reach a plateau, indicating that the agent has successfully optimized its control strategy. However, different configurations display varying convergence speeds. It is observed that employing multiple pairs of synthetic jets for flow control does not accelerate the reinforcement learning training process. The reward function for single-pair synthetic jets converges more rapidly than that for multiple pairs, suggesting that single-pair jet configurations are more conducive to faster convergence and more effective for the agent to learn the control strategy.

### v5

@@ 2.5 Enhanced Active Flow Control
  • 1.
    Agent : agent selection depends on the state space, action space, exploration-exploitation trade-off, computational resources, and objectives. For high-dimensional continuous spaces like fluid mechanics, PPO Rabault et al. (2019, 2020); Tang et al. (2020), Soft Actor-Critic (SAC) Wang et al. (2022); Jia and Xu (2024); Xia et al. (2024), Twin Delayed Deep Deterministic Policy Gradient (TD3) Fan et al. (2020) and Truncated Quantile Critics (TQC) Xia et al. (2024) are ideal, with PPO excelling in real-time flow control due to its exploration balance and efficiency.
  • 2.
@@
    Environment : the environment interacting with the agent is instantiated using CFD. Figure 6 illustrates the integration of CFD computations as a DRL training environment and demonstrates the benefits of accelerating the training process by using multiple parallel CFD environments. This parallelization strategy is inspired by the work of Rabault and Kuhnle (2019), and also incorporates insights from Jia and Xu (2024). The environment depends on the type of CFD solver, and there are various codes available for coupling numerical simulation environments with RL frameworks. RL has been integrated with the FEniCS environment Rabault et al. (2019), the OpenFOAM environment Wang et al. (2022), and the Nek5000 environment Li and Zhang (2022).
  • 3.
    Action $a_{t}$: the agent’s actions are the dimensionless mass flow rates of the synthetic jets, $Q_{i}$ ($i=1,2$), constrained by $Q_{1}+Q_{2}=0$ to ensure mass conservation and improve numerical stability. The control actions for the synthetic jets are defined such that a positive value indicates the jet is in the blowing state, while a negative value indicates the jet is in the suction state. To ensure the stability of the system and prevent non-physical large actuations that could destabilize it, a limit is imposed on the normalized mass flow rate of the jets. Additionally, to further enhance stability and mitigate any non-physical instabilities in the flow, a smoothing function is applied to the control actions.
  • 4.
    State $s_{t}$ : the agent observes the CFD environment by collecting physical field data from strategically placed probes. Inspired by Wang et al. (2024), Rabault et al. (2019), and Li and Zhang (2022), probes are positioned in wake regions with high instability and sensitivity, optimizing flow control by monitoring time-averaged fluctuating pressure. Additionally, multiple sensor probes are distributed around the cylinder to capture detailed physical information. Paris et al., Li and Zhang and Wang et al. studied the impact of the distribution of observation probes on control performance.
  • 5.
    Reward $r_{t}$: the reward function is designed to optimize flow control by reducing both drag and lift forces around the cylinder. To facilitate comparison with prior research, our reward function does not incorporate external energy consumption Rabault et al. (2019); Jia and Xu (2024); Wang et al. (2022); Jia and Xu (2024). We aim to train the agent to develop a low-energy control strategy through its interaction with the environment. It is formulated as:
@@
    where $C_{D,0}$ represents the baseline drag coefficient, providing a reference for potential reward improvement. The $\left(\cdot\right)_{T_{i}}$ represents the sliding average over the duration of one jet flow control period ${T_{i}}$. The parameter $\omega$ is a weighting factor that balances the importance of reducing lift relative to drag, ensuring that the reward function effectively incentivizes the agent to minimize both forces simultaneously.
  • 6.
    Interaction: the agent interacts with the environment at each time step $t$ by observing the current flow state $s_{t}$, selecting an action $a_{t}$ based on the policy $\pi_{\theta}(a_{t}\mid s_{t})$, and receiving a reward $r_{t}$ and the next state $s_{t+1}$. The agent iteratively refines $\pi_{\theta}$ to maximize cumulative rewards $\sum_{t}\gamma^{t}r_{t}$. The DRL training process encompasses 3,000 episodes, each comprising 100 timesteps. Each timestep corresponds to 0.025 non-dimensional time units, encompassing approximately 8 to 10 vortex shedding cycles. The PPO algorithm iteratively updates the policy using mini-batch optimization across all 3,000 episodes.
@@ 3 Results: flow control performance of a circular cylinder
This section examines the impact of synthetic jet positioning on flow control performance and costs, quantified by dimensionless mass flow rates. Five single-pair jet configurations, two two-pairs configurations, and one additional single-pair configuration are analyzed. It is important to note that during the training process, only the positions or quantities of the synthetic jets are altered, while the hyperparameters for each DRL training remain consistent.
In our DRL framework, a single agent is designed to control multiple actions corresponding to the manipulation of multiple synthetic jets. This unified control approach ensures that the agent optimizes its policy considering the overall system dynamics, rather than relying on cooperation among multiple agents. The agent receives a global state representation and outputs actions that collectively influence the environment, allowing for efficient and coordinated control of all jets. Importantly, the agent operates based on a single reward function that evaluates the overall performance of the system, ensuring that the control strategy is optimized for the collective impact of all jets.
Figure 7: Evaluation of flow control performance for various jet configurations under the condition of activating only a single pair of synthetic jets. (a) Reward curves over training episodes. (b) The time history of the $C_{D}$ of the cylinder in both baseline and controlled flow conditions. (c) The time history of the $C_{L}$ of the cylinder in both baseline and controlled flow conditions.
@@ 3.1.1 Scenario A: Control Performance of Jet Pair
In this section, we focus solely on the sensitivity of DRL training control performance to the angular positions of single pairs of synthetic jets on the cylinder surface. Specifically, we compare and analyze the reward function curves, flow control strategies, and flow control performance for synthetic jets at different positions. The range of angular positions on the cylinder surface spans from 60° to 120°. For instance, the jet configuration $J_{1}$ comprises a pair of synthetic jets, $J_{1}^{1}$ and $J_{1}^{2}$, with $J_{1}^{1}$ located at an angle of 120°. Similarly, the $J_{5}$ configuration includes two jets, $J_{5}^{1}$ and $J_{5}^{2}$, with $J_{5}^{1}$ positioned at 60°. The positions of the remaining synthetic jets are designed in a similar manner.
The reward function curves for different jet positions are illustrated in Figure 7 when only a single pair of synthetic jets is activated. These curves reflect the performance of the DRL agent during training. The initial fluctuations indicate the exploration phase, while the subsequent upward trend signifies the gradual optimization of the control strategy. Comparing the reward function curves for different jet configurations, we find that the jets located at position $J_{1}$ show the slowest growth and the lowest final reward, highlighting the challenges in control. The jets at position $J_{5}$ demonstrate moderate convergence performance, whereas the jets at positions $J_{2}$, $J_{3}$, and $J_{4}$ exhibit rapid convergence. After 1500 episodes, the reward curves stabilize, indicating that the DRL agent has found a stable and effective control strategy capable of robustly reducing $C_{D}$ and controlling $C_{L}$.
To clearly demonstrate the control performance based on DRL, the drag and lift coefficients for both the baseline and controlled flows are presented in figures 7(b) and 7(c). The baseline flow, devoid of active flow control, serves as a reference. In stark contrast, the controlled flow, modulated by the DRL algorithm, consistently achieves an 8% reduction in drag coefficient across all jet configurations. This reduction is not merely transient but is sustained over time, as evidenced by the stabilization of the drag and lift coefficients to a steady state following the initial transient phase. This suppression of periodic fluctuations underscores the robustness and effectiveness of the DRL-driven control mechanism in mitigating flow-induced instabilities.
@@
The velocity fields of the controlled flow for synthetic jets positioned at various angular orientations are shown in figure 8, highlighting the sensitivity of control performance to jet placement. Specifically, when the jets are positioned at 60°, the controlled flow still exhibits periodic vortex shedding, indicating that this orientation is less effective in stabilizing the flow. Conversely, when the jets are positioned at 120°, periodic vortex shedding is significantly suppressed, although some residual instabilities remain. This suggests that while the 120° orientation is more effective than 60°, it is not optimal. The most effective jet placements are at 75°, 90°, and 105°, where the wake of the cylinder flow is fully stabilized, and vortex shedding is completely eliminated. These results underscore the critical role of jet orientation in achieving optimal flow control and highlight the potential for further optimization through precise placement of control actuators.
@@ 3.1.2 Scenario A: Energy Consumption of Jet Pair
Figure 9: Energy consumption analysis of jet pairs positioned at different angular locations. (a) The control action values over non-dimensional time for each jet configuration. The integral of the control action values over time represents the total external energy consumed by the flow control system. (b) The the angular placement of jet pairs $J_{1}$ through $J_{5}$ around the cylinder.
The relationship between the control action values of each jet configuration and the dimensionless time is depicted in figure 9, offering insights into the energy consumption required for the desired flow control. It also shows the azimuthal positions of five pairs of injectors along the surface of the cylinder. Among them, the jet position $J_{1}$ is located closest to the downstream stagnation point, while the jet position $J_{5}$ is situated closer to the upstream side. The control action value of the synthetic jet represents the intensity of the jet-driven force, and the integral of the control action value over time signifies the total external energy consumed to implement flow control.
As depicted in figure 9, at the initial stage of flow control, the control effect curves of each jet configuration exhibit a peak, indicating the maximum energy required at the onset of jet actuation. Following this initial transient phase, the action values stabilize, suggesting that the control intensity and energy consumption become more stable and energy-efficient over time. Notably, the curves for $J_{5}$ and $J_{1}$ display the highest initial action values, indicating that these positions require more initial energy to initiate flow control compared to other configurations. In contrast, the action values for the synthetic jets positioned at $J_{2}$, $J_{3}$, and $J_{4}$ stabilize at a lower level than those of $J_{5}$ and $J_{1}$, demonstrating that these configurations are more energy-efficient in maintaining flow control.
In the later stages of flow control, the synthetic jet at $J_{2}$ operates effectively with only 1% of the incoming flow rate, while those at $J_{1}$ and $J_{5}$ require over 3% to achieve comparable control performance. This disparity suggests that injectors positioned at 60° and 120° exhibit lower energy efficiency, necessitating a higher input to sustain the desired control effect. In contrast, injectors at 90° and 105° demonstrate superior efficiency, maintaining effective flow control with significantly lower energy consumption. Such findings highlight the importance of selecting optimal jet positions to minimize energy expenditure while achieving targeted flow control objectives. This insight highlights the critical importance of optimizing jet positioning to minimize energy consumption, a key consideration in designing efficient flow control systems that maximize aerodynamic performance while reducing power requirements around the cylinder.
@@ 3.2.1 Scenario B: Control Performance of Multiple Jets Control
Based on the results of flow control using a single pair of synthetic jets, it is observed that the configurations with jets positioned at 60° and 120°, namely $J_{5}$ and $J_{1}$, consume the most energy. Consequently, when employing multiple pairs of synthetic jets for flow control, the jets positioned at 60° and 120° are excluded. Instead, combinations involving $J_{2}$, $J_{3}$, and $J_{4}$ are explored. The tested configurations include two pairs of synthetic jets, specifically $J_{2}$ and $J_{3}$ as well as $J_{3}$ and $J_{4}$, and a configuration with three pairs, namely $J_{2}$, $J_{3}$, and $J_{4}$.
Figure 10: Control performance evaluation of multiple jet configurations around a cylinder. (a) Reward curves over training episodes for various multi-jet configurations. (b) $C_{D}$ over non-dimensional time. (c) $C_{L}$ over non-dimensional time.
In figure 10, we present the control performance and convergence rates for flow control using multiple pairs of synthetic jets. To provide a comprehensive comparison, we also include the results of three single-pair synthetic jet controls, thereby contrasting the performance of multiple jet combinations with that of individual pairs. The reward curves in the figure reflect the effectiveness of the DRL agent in learning flow control strategies using different synthetic jet configurations. Specifically, figure 10(a) offers a detailed analysis of the convergence performance during the training of deep reinforcement learning. In all configurations, the reward function exhibits an increasing trend during the initial training phase. As the DRL training progresses, the reward values gradually stabilize, indicating that the agent has successfully optimized the control strategy. However, different jet configurations demonstrate varying convergence rates and maximum reward values. Interestingly, we observe that employing multiple pairs of synthetic jets simultaneously for flow control does not accelerate the reinforcement learning training process. In fact, the reward functions for single-pair synthetic jets converge more rapidly than those for multiple pairs. This suggests that single-pair jet configurations are more conducive to faster convergence.